OpenStudy (anonymous):

Gino designed a garden for his rectangular backyard. The design he drew on paper is 6 inches wide. The width of his backyard is thirty feet. If his drawing is to scale, and his backyard is forty feet long, how many inches long is his drawing? Write a proportion using words. Then, use the proportion to calculate the drawing length. Show your work.

OpenStudy (campbell_st):

scale for the width 6 inches = 30 ft or 1 inch = 5 feet 1:60 same for the length then 8 inches = 40 ft or 1 inch = 5 ft

OpenStudy (campbell_st):

the calculation is \[\frac{1}{5}=\frac{x}{40}\] \[x = \frac{ 40 \times1}{5}\]

OpenStudy (anonymous):

??? dont u use cross mutiplication

OpenStudy (anonymous):

answer is 8